The table below shows the probability distribution for x, the number that shows up when a loaded die is rolled:

X P(X)
1 0.12
2 0.10
3 0.12
4 0.14
5 0.15
6 0.37

a) Find the mean of the distribution.

b) Find the standard deviation of the distribution.
Answer #1

Solution:

x P(x) x * P(x) x2 * P(x)
1 0.12 0.12 0.12
2 0.1 0.2 0.4
3 0.12 0.36 1.08
4 0.14 0.56 2.24
5 0.15 0.75 3.75
6 0.37 2.22 13.32
Sum 1 4.21 20.91

Mean = X * P(X)

= 4.21

Standard deviation =

=X 2 * P(X) - 2

=  20.91 - (4.21)2 = 3.1859

Standard deviation = = 1.7849